OpenStudy (anonymous):

A high school purchases a printing machine that can put lettering on clothing for $6000. The school plans to print T-shirts with the school logo and sell them. Each T-shirt costs the school $4 and will sell for $12. How many T-shirts must the school sell before breaking even?

OpenStudy (ganpat):

Break even point is when total expense is equal to total income. total cost for machine is 6000$ each t shirt will be costing another 4$ and they plan to sell it for 12$ let x be total t shirt So.. for Break Even point: 12x = 6000+4x calculate for x

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

well I tried doing it but the answers to the question in 700 T-shirts, 750 T-shirts, 1500 T-shirts, and 7500 T-shirts.

OpenStudy (ganpat):

you tried and what is your answer ?

OpenStudy (anonymous):

375

OpenStudy (ganpat):

12x = 6000 + 4x 8x = 6000 x = 750 T shirts...

OpenStudy (ganpat):

i mean.. my calculation says.. if you could write what you did, i can correct it...

OpenStudy (anonymous):

oh... sorry I feel stupid but I get it now! Thanks!!!